News about Science Data Product ForumWe would like to draw your attention about the existence of the Science Data Products internet forum which can be found at: http://www.eso.org/sci/data-processing/forum.html
ISAAC NewsISAAC is being offered again but will be decommissioned during Period 92 to allow the installation of SPHERE. The forseen date for the start of the decommissioning is 10th December 2013.
- Upper quadrant of HAWAII detector is off: We restrict observations to small objects or point sources where nodding on only the lower half of the detector is sufficient.
- M7 in simulation: M7, the mirror to select between SW and LW arm is not working properly and for the moment is set to SW arm. We plan to change to LW again in November to allow the observation of P92 LW programs.
- Grating reproducibility: We have problems with the reproducibility of the grating position. If the central wavelength is changed, offsets up to 20 pixels are observed. This can be a problem when more then one setup is used in an OB and the telluric standard is taken afterwards after the grating was moved. For the time being, we recommend that OBs contain not more than two setups, and if they contain two setups, clearly mark this in the README file. In those cases, we will try to take a telluric before with the first setup and one afterwards with the second setup, so that no shifts occur between the science target and the telluric.
- More information can be found on the QC problem page.
- 30th July 2013 The M7 was moved to oeprate ISAAC in SW mode from now on.
- 28th June 2013 The M7 was partly fixed during the intervention but it is not clear how stable the current solution is. We will therefore limit the movement of the M7 to the absolute minimum. For the moment, we are observing in LW arm.
- 27th June 2013 The upper left quadrant of the Hawaii detector is still bad. We will not be able to fix it before the decommissioning of ISAAC. For the observations, we move the targets to use the lower half of the detector only. This limits the field of view to 2.5x1.25 arcmin, larger nodthrows are not recommended.
- 13th June 2013 New intervention to fix the problem with the M7 as well as the HAWAII detector planned for June 16 - 26.
- 26th May 2013 The 2nd intervention first seemed to have been successfull. However, after a few days, the problem with M7 re-appeared, and also the upper left quadrant of the SW-detector (HAWAII) is off.
- 6th May 2013 1st intervention tp fix M7 was unsuccessful. New intervention will probably start on 14th May.
- 6th March 2013: Hawaii read speed factor changed to 8 because of reappearance of OE effect.
- 23rd July 2012: Hawaii read speed factor changed to 6 because OE effect is gone.
- 24th June 2012: Hawaii read speed factor changed to 10 because of strong OE effect.
- 24th February 2012: ISAAC will likely be decommissioned during Period 90 to allow the installation of SPHERE.
- 18th November 2011: Short intervention on warm parts of ISAAC (electronics) successful. Several reproducibility problems could be solved.
- 12th November 2011: Hawaii read speed factor changed to 4 because OE effect is gone.
- 30th September 2011: Hawaii read speed factor changed to 11 to cope with increasing OE effect.
- 14th September 2011: Odd Even effect is back on Hawaii detector, read speed factor changed to 8.
- 14th August 2011: Read speed of Hawaii detector changed back.
- July 2011: The detector intervention was successfull, the old HAWAII detector could be restored. ISAAC has been successfully recovered and is in full operation again.
- May 2011:On 2011-05-17T12:18 during the daytime calibrations the upper left quadrant of the Hawaii detector died. The detector was disabled, thus the SW arm is currently not available. An intervention is planned for June to either fix the detector or replace it.
- January 2011: Emergency intervention on the autoguider. Instrument was detached from the telescope but successfully recovered on 25th January.
- February 2010: Emergency intervention to fix the filter wheel 2. ISAAC was successfully recovered on 26th February.
- P84: In August 2009, ISAAC was moved to the Nasmyth A focus on UT3 (Melipal). Manuals and CfP for P84 have been updated accordingly.
This document was last updated: 2013-07-30